# Functional Regression with Nonstationarity and Error Contamination: Application to the Economic Impact of Climate Change

### v1

@@ Abstract
This paper studies a functional regression model with nonstationary dependent and explanatory functional observations, in which the nonstationary stochastic trends of the dependent variable are explained by those of the explanatory variable, and the functional observations may be error-contaminated. We develop novel autocovariance-based estimation and inference methods for this model. The methodology is broadly applicable to economic and statistical functional time series with nonstationary dynamics. To illustrate our methodology and its usefulness, we apply it to the evaluation of the global economic impact of climate change, an issue of intrinsic importance.
@@ 1 Introduction
Only recently has the literature begun to consider nonstationary dependent observations, even if many economic and statistical functional time series tend to be nonstationary, as noted in recent papers (e.g., Chang et al., 2016b, ; Beare et al.,, 2017; Franchi and Paruolo,, 2020; Li et al.,, 2023; Nielsen et al.,, 2023, 2024; Seo,, 2024; Seo and Shang,, 2024). As a result, statistical methods developed for such time series are currently limited to analyzing their essential properties, such as cointegration, stochastic trends, and the dominant subspace. Despite its empirical relevance, articles aiming to develop inferential methods for functional (auto-)regression models involving nonstationary functional time series are scarce; to the best of the authors’ knowledge, there are currently only a few preprints of papers (e.g., Chang et al., 2016a, ; Chang et al.,, 2024). We fill this gap by developing novel statistical methods for functional regression models where both regressand and regressor exhibit unit-root-type nonstationary behavior allowing cointegration, a feature particularly important for economic and financial applications.
@@
In addition to incorporating nonstationarity into functional regression models, we aim to enhance the real-world applicability of our methods by addressing a typical and practical aspect of functional data that has recently been discussed in the literature: incomplete and partially observed data (see, e.g., Chen et al.,, 2022; Seong and Seo,, 2025). In the majority of real data analyses, (i) each functional observation, say $x_{t}(u)$ for $u\in[a_{1},a_{2}]$, is not directly observed, and often constructed by its partial and discrete realizations $(x_{t}(u_{1}),\ldots,x_{t}(u_{n}))^{\prime}$ for $u_{1},\ldots,u_{n}\in[a_{1},a_{2}]$ and (ii) oftentimes, the number of discrete observations $n$ is not large enough. In fact, (i) and (ii) are pointed out by Seong and Seo, (2025) in the context of functional linear models, and they argued that “endogeneity” caused by measurement errors need to be properly addressed for estimation and inference (see also Chen et al.,, 2022). As a specific example, consider a case where functional observations are probability density-valued (as in Section 5 to appear). This case has gained significant interest in the literature; for the stationary case, see e.g., Kneip and Utikal, (2001) and Park and Qian, (2012), while for the nonstationary case refer to Chang et al., 2016b and Seo and Beare, (2019). In this scenario, the true probability density is not observable, and thus, it needs to be replaced by a proper nonparametric estimate. This naturally introduces small or large measurement errors in practice. In this paper, we explicitly consider cases where the variables of interest, which are nonstationary, are also error-contaminated, and then pursue statistical methods that are robust to error contamination. This not only distinguishes the present paper significantly from existing works (cf., e.g., Benatia et al.,, 2017; Park and Qian,, 2012; Chen et al.,, 2022; Babii,, 2022; Seong and Seo,, 2025) but also makes our proposed methods more appealing to applied researchers. We also believe that our methodology can be applied to various economic and financial time series.
@@
More technically and specifically, we assume that the variables of interest are cointegrated functional time series, following the framework of Chang et al., 2016b and Beare et al., (2017). This assumption has been widely used in the recent literature on nonstationary functional time series, especially in economic applications (see, e.g., Nielsen et al.,, 2023; Seo,, 2024). We then assume that these variables can only be observed with additive measurement errors. As noted by Seong and Seo, (2025), the problem of neglected error contamination generally results in the inconsistency of standard estimators constructed from the sample covariance operator $\widehat{C}_{0}$ of the regressor (this is also true in our model, which will be discussed in Section 5 in more detail). This inconsistency arises primarily because $\widehat{C}_{0}$ is inherently contaminated by measurement errors and, consequently, becomes a distorted estimator of its population counterpart. To address this issue of error contamination, we consider autocovariance-based inference, avoiding the direct use of the covariance operator of an error-contaminated variable for statistical inference, as in some recent articles on functional regression models (see, e.g., Chen et al.,, 2022). More specifically, we construct our proposed estimator based on the lag-$\kappa$ sample autocovariance $\widehat{C}_{\kappa}$ for some positive $\kappa$. This approach is grounded in the observation that, as long as the measurement errors are not strongly correlated and satisfy certain mild regularity conditions (to be detailed), (i) the sample autocovariance $\widehat{C}_{\kappa}$ will be less affected by measurement errors, and (ii) the assumption that measurement errors are not strongly correlated does not seem overly restrictive, given that such errors in functional data analysis commonly arise from constructing each individual functional observation based on its discrete realizations; as will be detailed in Section 5, our asymptotic analysis requires a much weaker condition on the serial correlation of the measurement errors rather than complete serial uncorrelatedness. It should be noted that, in this paper, we also consider the case $\kappa=0$, which yields the standard covariance-based estimator in the functional linear model, and study its detailed asymptotic properties, a contribution that is, to the best of the authors’ knowledge, also novel.
@@
We develop relevant autocovariance-based inferential methods that are robust to the potential presence of measurement errors. This includes a novel dimension-reduction method, our proposed estimator of the slope parameter in the functional regression model based on it, and their asymptotic properties. The proposed estimator, to some extent, resembles the conventional two-step estimator of Engle and Granger, (1987) in that both use residuals computed from the estimated relationship between the nonstationary components of the model; however, beyond this, the two approaches differ substantially in structure and purpose. We also provide numerical studies with real-world data and simulation experiments to examine the finite-sample properties of our proposed estimator. As an application, we illustrate the empirical relevance of our proposed methodology by considering the empirical model for studying the global economic impact of climate change. Specifically, we show that the proposed framework effectively estimates the distributional relationship between land temperature anomalies, often considered a measure of climate change, and regional economic growth rates under the possible presence of measurement errors, thereby offering a robust basis for assessing heterogeneous climate–economy relationships across the globe.
@@
The rest of the paper is organized as follows. Section 2 briefly reviews essential preliminaries on nonstationary cointegrated functional time series. Section 3 describes the model considered in this paper, and Section 4 develops the inferential methods for the model. In Section 5, we apply the proposed method to examine the global economic impact of climate change. Section 6 concludes.
@@ 2.1 Notation and simplification
Section A of the Appendix reviews basic concepts on bounded linear operators and random elements associated with two (possibly different) Hilbert spaces. Accordingly, we let $\mathcal{L}_{\mathcal{H}}$ denote the space of bounded linear operators on $\mathcal{H}$ with the usual operator norm $\|\cdot\|_{\textrm{op}}$, and let $\otimes$ denote the tensor product associated with $\mathcal{H}$, $\mathcal{H}_{y}$, or both (see (A.26)). Section A also reviews $\mathcal{H}$-valued random elements $X$, their expectation (denoted $\mathbb{E}[X]$), covariance operator (denoted $C_{X}:=\mathbb{E}[(X-\mathbb{E}(X))\otimes(X-\mathbb{E}(X))]$), and the cross-covariance with an $\mathcal{H}_{y}$-valued random element $Y$ (denoted $C_{XY}:=\mathbb{E}[(X-\mathbb{E}(X))\otimes(Y-\mathbb{E}(Y))]$). In addition, for $A\in\mathcal{L}_{\mathcal{H}}$, concepts such as the adjoint (denoted $A^{\ast}$), range (denoted $\operatorname{ran}A$), and kernel (denoted $\ker A$), as well as properties such as self-adjointness, compactness, Hilbert–Schmidtness, and nonnegativity are introduced in that section, and they will be useful in the subsequent discussion.
We will consider sequences of random linear operators, constructed from random elements in $\mathcal{H}$ and $\mathcal{H}_{y}$ (for a more detailed discussion on general random linear operators, see Skorohod,, 1983). For any such operator-valued random sequence $\{A_{j}\}_{j\geq 1}$, we write $A_{j}\to_{p}A$ to denote convergence in probability with respect to the operator norm (i.e., $\|A_{j}-A\|_{\textrm{op}}\to_{p}0$). In the subsequent discussion, convergence in probability sometimes occurs for $\mathcal{H}$- or $\mathcal{H}_{y}$-valued elements (in the appropriate norm), but for convenience we use the same notation $\to_{p}$ to denote such convergence throughout, as distinguishing between the two would add notational complexity with little benefit. Moreover, as is common in the literature (see e.g., Seo,, 2024), we write $A_{j}=A+O_{p}(a_{T})$ (resp. $A_{j}=A+o_{p}(a_{T})$) if $\|A_{j}-A\|_{\textrm{op}}=O_{p}(a_{T})$ (resp. $\|A_{j}-A\|_{\textrm{op}}=o_{p}(a_{T})$) for some sequence $a_{T}$. For any two operators $A$ and $B$, we write $A=_{d}B$ to denote equivalence in their finite dimensional distributions as in Seo, (2024), i.e., $A=_{d}B$ if for any $n>0$, $\{v_{j}\}_{j=1}^{n}$ ( $\subset\mathcal{H}\text{ or }\mathcal{H}_{y}$) and $\{w_{j}\}_{j=1}^{n}$ ( $\subset\mathcal{H}\text{ or }\mathcal{H}_{y}$), the distribution of $(\langle Av_{1},w_{1}\rangle,\ldots,\langle Av_{n},w_{n}\rangle)^{\prime}$ equals that of $(\langle Bv_{1},w_{1}\rangle,\ldots,\langle Bv_{n},w_{n}\rangle)^{\prime}$.
@@ 2.2 Cointegrated $\mathcal{H}$-valued time series
Subsequently, we will consider the cointegrated time series introduced in this section, but some additional restrictions will be imposed for our asymptotic analysis in Section 3.
@@ 3 Proposed model
Let $\{x_{t}\}_{t\geq 1}$ be a cointegrated $\mathcal{H}$-valued time series, as detailed in Section 2.2, which induces a bipartite partition of $\mathcal{H}$ into a nonstationary subspace $\mathcal{H}^{N}$ and a stationary subspace $\mathcal{H}^{S}$. We consider the following data generating mechanism:
$y_{t}=f(x_{t})+u_{t},\quad f:\mathcal{H}\to{\mathcal{H}}_{y},$ (2)
$u_{t}^{N}={\rm P}^{N}\Delta x_{t}=\sum_{j=0}^{\infty}\psi_{j}^{N}\epsilon_{t-j},\quad u_{t}^{S}={\rm P}^{S}x_{t}=\sum_{j=0}^{\infty}\psi_{j}^{S}\epsilon_{t-j}.$ (3)
@@
Throughout this paper, we assume that $x_{t}$ cannot be directly observed but that only $\tilde{x}_{t}$, observed with measurement errors, is available. As highlighted by Seong and Seo, (2025), this assumption is empirically relevant because functional observations used in practice are often incompletely observed, with only finitely many discrete realizations available to practitioners. Consequently, it is common to construct a functional observation $z_{t}$ in advance by smoothing its $n$ discrete data points $z_{t}(s_{1}),\ldots,z_{t}(s_{n})$, with $s_{j}$ included in the entire interval $a_{1},a_{2}]$, before computing estimators or test statistics. While one may disregard measurement errors for simplicity if $n$ is large enough and the data points are densely observed over $a_{1},a_{2}]$, this is often not the case in practice (our empirical application in Section 5 is an example). We develop the theoretical results under the presence of measurement errors in functional variables, while also discussing how these results simplify in the absence of such errors. Accordingly, the subsequent theoretical developments remain applicable to the error-free case, which has more commonly been considered in the functional data analysis literature.
@@
where $e_{t}$ may generally be correlated with the variables $u_{t}^{N}$ and $u_{t}^{S}$ appearing (2) and (3), and it may also be serially correlated. A key assumption, which we employ for our asymptotic analysis, is that $e_{t-\kappa}$ (and also $e_{t+\kappa}$) for some finite $\kappa>0$ has asymptotically negligible sample (cross-)covariance with $u_{t}^{N}$, $u_{t}^{S}$ and $e_{t}$. Given that $e_{t-\kappa}$ or $e_{t+\kappa}$ is (mostly) smoothing error or random disturbance associated with a variable observed at a different time, this assumption is practically reasonable and more likely to be satisfied even for a small positive $\kappa$. Of course, $y_{t}$ can also suffer from similar error contamination but, as in the conventional multivariate regression model, its measurement error is absorbed into $u_{t}$. This only changes the interpretation of $u_{t}$ in the subsequent analysis. Under the presence of measurement errors $e_{t}$, we may rewrite (2) as follows:
@@
    (Error-contaminated case) $\kappa\geq 1$ and $\mathbb{E}[e_{t}\otimes z_{t+\ell}]=0$ any $|\ell|\geq\kappa$, where $z_{t}=e_{t}$, $u_{t}^{N}$ and $u_{t}^{S}$; moreover, $\{e_{t}\}_{t\geq 1}$ is stationary and geometrically strongly mixing, $\mathbb{E}[\|e_{t}\|^{4}]<\infty$, $T^{-1}\sum_{t=1}^{T}e_{t}=O_{p}(T^{-1/2})$, and $T^{-1}\sum_{t=1}^{T}e_{t}\otimes z_{t+\ell}=\mathbb{E}[e_{t}\otimes z_{t+\ell}]+O_{p}(T^{-1/2})$ for any $|\ell|\geq\kappa$.
@@
Some comments on Assumptions 1 and E are in order. We assume that $\mathcal{H}_{y}=\mathbb{R}$ (resp. $\mathcal{H}_{y}=\mathcal{H}$) if $y_{t}$ is scalar-valued (resp. function-valued). In the function-valued case, $y_{t}$ may be defined not on $a_{1},a_{2}]$, as $x_{t}$ is, but on a different interval, say $b_{1},b_{2}]$; extending or applying the subsequent theoretical results to this case is straightforward, and assuming $\mathcal{H}_{y}=\mathcal{H}$ entails no loss of generality. In Assumption 1(a), we assume that $d_{N}$ is finite. This condition has been widely employed in the literature on nonstationary functional time series and also seems empirically relevant (Chang et al., 2016b, ; Nielsen et al.,, 2023; Seo and Shang,, 2024). Moreover, a wide class of functional time series satisfies this condition (see Remark 3.2). For convenience, we also assume that $d_{N}$ is known, even though it is unknown to practitioners in most empirical applications. However, replacing $d_{N}$ with various consistent estimators does not affect the asymptotic results to be developed. Moreover, we show in Section C.2 that the variance-ratio testing procedure of Nielsen et al., (2023) can be used in our setting, allowing for the presence of measurement errors. Assumption 1(b) is employed to facilitate our theoretical analysis based on useful limit theorems in the existing literature (see, e.g., Bosq,, 2000). Given Assumption 1(b), Assumption 1(c) does not appear restrictive, and some primitive sufficient conditions can be found in (Bosq,, 2000, Chapter 2). Assumption 1(d) is a technical condition required for our asymptotic analysis. Similar assumptions have been employed by Seo, (2024) in the study of FPCA for cointegrated time series in a Hilbert space setting. Moreover, sufficient, but not restrictive, conditions for the weak convergence results stated in Assumption 1(d) can be found in e.g., Berkes et al., (2013) and Seo, (2024).
@@
Assumption E states requirements on the measurement errors. Although our primary focus is on the error-contaminated case (Assumption E(a)), we will also show how the theoretical results simplify in the error-free case (Assumption E(b)) when applying the standard covariance-based approach (see Remark 3.1). If $e_{t}$ is serially independent and also independent of $u_{s}^{S}$ and $u_{s}^{N}$ for every $s$ and $t$, then, noting that (i) $\mathbb{E}[e_{t}\otimes u_{s}^{N}]=\mathbb{E}[e_{t}\otimes u_{s}^{S}]=0$ for all $s$ and $t$ in the considered scenario and (ii) $e_{t}\otimes z_{t}$ is a Hilbert-valued random variable (see Theorems 2.7 and 2.16 of Bosq,, 2000), the conditions in Assumption E(a) are satisfied under mild assumptions. However, Assumption E(a) is not restricted to such a case and allows more general cases; we specifically, $e_{t}$ is assumed to be uncorrelated with $z_{t+\ell}$ if $|\ell|$ is sufficiently large, while no restriction is imposed on $\mathbb{E}[e_{t}\otimes z_{t+\ell}]$ if $|\ell|$ is small. That is, for our theoretical investigation of the proposed method, we only require each of $e_{t}$, $u_{t}^{N}$, and $u_{t}^{S}$ to be uncorrelated with a non-adjacent past or future measurement error $e_{s}$. This not only seems to be a mild assumption but also reasonable for most empirical applications.
@@
It may be of interest to practitioners to examine the theoretical results for the standard FPCA-based estimator (corresponding to the case with $\kappa=0$, as will be shown) in the absence of measurement errors, since functional data may sometimes be observed accurately. To the authors’ knowledge, even in this simplified setting, no statistical theory has been established for nonstationary $y_{t}$ and $x_{t}$ (although the nonstationary functional AR(1) model was studied by Chang et al.,, 2024). Accordingly, the subsequent results for $\kappa=0$ and the error-free case are also novel, motivating our explicit consideration of this scenario.
@@
For the subsequent discussion, it is convenient to introduce additional notation. Whenever these quantities are well-defined, let $\bm{\mathsf{\lambda}}_{j}[A]$ as the $j$-th largest eigenvalue of a compact operator $A$, $\bm{\mathsf{v}}_{j}[A]$ as the corresponding eigenvector, and $\bm{{\Pi}}_{j}[A]$ as the orthogonal projection onto the span of $\bm{\mathsf{v}}_{j}[A]$; that is,

### v3

@@ Abstract
This paper studies a regression model with functional dependent and explanatory variables, both of which exhibit nonstationary dynamics. The model assumes that the nonstationary stochastic trends of the dependent variable are explained by those of the explanatory variables, and hence that there exists a stable long-run relationship between the two variables despite their nonstationary behavior. We also assume that the functional observations may be error-contaminated. We develop novel autocovariance-based estimation and inference methods for this model. The methodology is broadly applicable to economic and statistical functional time series with nonstationary dynamics. To illustrate our methodology and its usefulness, we apply it to evaluating the global economic impact of climate change, an issue of intrinsic importance.
@@ 1 Introduction
Only recently has the literature begun to consider nonstationary dependent observations, even if many economic and statistical functional time series tend to be nonstationary, as noted in recent papers (e.g., Chang et al., 2016b, ; Beare et al.,, 2017; Franchi and Paruolo,, 2020; Li et al.,, 2023; Nielsen et al.,, 2023, 2024; Seo,, 2024; Seo and Shang,, 2024). As a result, statistical methods developed for such time series are currently limited to analyzing their essential properties, such as cointegration, stochastic trends, and the dominant subspace. Despite its empirical relevance, articles developing inferential methods for functional (auto-)regression models involving nonstationary functional time series are scarce; to the best of the authors’ knowledge, there are currently only a few preprints (e.g., Chang et al., 2016a, ; Chang et al.,, 2024). We fill this gap by developing novel statistical methods for functional regression models where both regressand and regressor exhibit unit-root-type nonstationary behavior allowing cointegration, a feature particularly important for economic and financial applications.
@@
In addition to incorporating nonstationarity into functional regression models, we aim to enhance the real-world applicability of our methods by addressing a typical and practical aspect of functional data that has recently been discussed in the literature: incomplete and partially observed data (see, e.g., Chen et al.,, 2022; Seong and Seo,, 2025). In the majority of real data analyses, (i) each functional observation, say $x_{t}(u)$ for $u\in[a_{1},a_{2}]$, is not directly observed, and often constructed from its partial and discrete realizations $(x_{t}(u_{1}),\ldots,x_{t}(u_{n}))^{\prime}$ for $u_{1},\ldots,u_{n}\in[a_{1},a_{2}]$ and (ii) the number of discrete observations $n$ is often not large enough. In fact, (i) and (ii) are pointed out by Seong and Seo, (2025) in the context of functional linear models, and they argued that “endogeneity” caused by measurement errors need to be properly addressed for estimation and inference (see also Chen et al.,, 2022). As a specific example, consider a case where functional observations are probability density-valued (as in Section 5 to appear). This case has gained significant interest in the literature; for the stationary case, see e.g., Kneip and Utikal, (2001) and Park and Qian, (2012), while for the nonstationary case refer to Chang et al., 2016b and Seo and Beare, (2019). In this scenario, the true probability density is not observable, and thus, it needs to be replaced by a proper nonparametric estimate. This naturally introduces small or large measurement errors in practice. In this paper, we explicitly consider cases where the variables of interest, which are nonstationary, are also error-contaminated, and then develop statistical methods that are robust to error contamination. This not only distinguishes the present paper significantly from existing works (cf., e.g., Benatia et al.,, 2017; Park and Qian,, 2012; Chen et al.,, 2022; Babii,, 2022; Seong and Seo,, 2025) but also makes our proposed methods more appealing to applied researchers. We also believe that our methodology can be applied to various economic and financial time series.
@@
More technically and specifically, we assume that the variables of interest are cointegrated functional time series, following the framework of Chang et al., 2016b and Beare et al., (2017). This assumption has been widely used in the recent literature on nonstationary functional time series, especially in economic applications (see, e.g., Nielsen et al.,, 2023; Seo,, 2024). We then assume that these variables can only be observed with additive measurement errors. As noted by Seong and Seo, (2025), the problem of neglected error contamination generally results in inconsistency of standard estimators constructed from the sample covariance operator $\widehat{C}_{0}$ of the regressor (this is also true in our model, as will be discussed in Section 5 in greater detail). This inconsistency arises primarily because $\widehat{C}_{0}$ is inherently contaminated by measurement errors and, consequently, becomes a distorted estimator of its population counterpart. To address this issue of error contamination, we consider autocovariance-based inference, avoiding the direct use of the covariance operator of an error-contaminated variable for statistical inference, as in some recent articles on functional regression models (see, e.g., Chen et al.,, 2022). More specifically, we construct our proposed estimator based on the lag-$\kappa$ sample autocovariance $\widehat{C}_{\kappa}$ for some positive $\kappa$. This approach is grounded in the observation that, as long as the measurement errors are not strongly correlated and satisfy certain mild regularity conditions (to be detailed), (i) the sample autocovariance $\widehat{C}_{\kappa}$ will be less affected by measurement errors, and (ii) the assumption that measurement errors are not strongly correlated does not seem overly restrictive, given that such errors in functional data analysis commonly arise from constructing each individual functional observation based on its discrete realizations; as will be detailed in Section 5, our asymptotic analysis requires a much weaker condition on the serial correlation of the measurement errors rather than complete serial uncorrelatedness. It should be noted that, in this paper, we also consider the case $\kappa=0$, which yields the standard covariance-based estimator in the functional linear model, and study its detailed asymptotic properties, a contribution that is, to the best of the authors’ knowledge, also novel.
@@
We develop autocovariance-based inferential methods that are robust to the potential presence of measurement errors. This includes a novel dimension-reduction method, our proposed estimator of the slope parameter in the functional regression model, and their asymptotic properties. The proposed estimator bears some resemblance to the conventional two-step estimator of Engle and Granger, (1987) in that both use residuals computed from the estimated relationship between the nonstationary components of the model; however, beyond this superficial similarity, the two approaches differ substantially in structure and purpose. We also provide numerical studies with real-world data and simulation experiments to examine the performance of our proposed estimator. As an application, we illustrate the empirical relevance of our proposed methodology by applying it to an empirical model for studying the global economic impact of climate change. Specifically, we show that the proposed framework effectively estimates the distributional relationship between land temperature anomalies–often considered a measure of climate change–and regional economic growth rates in the possible presence of measurement errors, thereby offering a robust basis for assessing heterogeneous climate–economy relationships across the globe.
@@
The rest of the paper is organized as follows. Section 2 reviews essential preliminaries on nonstationary cointegrated functional time series. Section 3 details the regression model, and Section 4 develops inferential methods for it. In Section 5, we apply the proposed method to examine the global economic impact of climate change. Section 6 concludes.
@@ 2.1 Notation and simplification
Section A of the Supplementary Material reviews basic concepts of bounded linear operators and random elements associated with two (possibly different) Hilbert spaces. Accordingly, we let $\mathcal{L}_{\mathcal{H}}$ denote the space of bounded linear operators on $\mathcal{H}$ with the usual operator norm $\|\cdot\|_{\textrm{op}}$, and let $\otimes$ denote the tensor product associated with $\mathcal{H}$, $\mathcal{H}_{y}$, or both (see (A.26)). Section A also reviews $\mathcal{H}$-valued random elements $X$, their expectation (denoted $\mathbb{E}[X]$), covariance operator (denoted $C_{X}:=\mathbb{E}[(X-\mathbb{E}(X))\otimes(X-\mathbb{E}(X))]$), and cross-covariance with an $\mathcal{H}_{y}$-valued random element $Y$ (denoted $C_{XY}:=\mathbb{E}[(X-\mathbb{E}(X))\otimes(Y-\mathbb{E}(Y))]$). For $A\in\mathcal{L}_{\mathcal{H}}$, concepts such as the adjoint (denoted $A^{\ast}$), range (denoted $\operatorname{ran}A$), and kernel (denoted $\ker A$), as well as properties such as self-adjointness, compactness, Hilbert–Schmidtness, and nonnegativity are introduced in that section, and they will be useful subsequently.
We will consider sequences of random linear operators, constructed from random elements in $\mathcal{H}$ and $\mathcal{H}_{y}$ (for a more detailed discussion on general random linear operators, see Skorohod,, 1983). For any such operator-valued random sequence $\{A_{j}\}_{j\geq 1}$, we write $A_{j}\to_{p}A$ to denote convergence in probability with respect to the operator norm (i.e., $\|A_{j}-A\|_{\textrm{op}}\to_{p}0$). In the subsequent discussion, convergence in probability sometimes occurs for $\mathcal{H}$- or $\mathcal{H}_{y}$-valued elements (in the appropriate norm), but for convenience we use the same notation $\to_{p}$ to denote such convergence throughout, as distinguishing between the two would add notational complexity with little benefit. Moreover, as is common in the literature (see, e.g., Seo,, 2024), we write $A_{j}=A+O_{p}(a_{T})$ (resp. $A_{j}=A+o_{p}(a_{T})$) if $\|A_{j}-A\|_{\textrm{op}}=O_{p}(a_{T})$ (resp. $\|A_{j}-A\|_{\textrm{op}}=o_{p}(a_{T})$) for some sequence $a_{T}$. For any two operators $A$ and $B$, we write $A=_{d}B$ to denote equivalence in their finite-dimensional distributions as in Seo, (2024), i.e., $A=_{d}B$ if, for any $n>0$, $\{v_{j}\}_{j=1}^{n}$ ( $\subset\mathcal{H}\text{ or }\mathcal{H}_{y}$) and $\{w_{j}\}_{j=1}^{n}$ ( $\subset\mathcal{H}\text{ or }\mathcal{H}_{y}$), the distribution of $(\langle Av_{1},w_{1}\rangle,\ldots,\langle Av_{n},w_{n}\rangle)^{\prime}$ equals that of $(\langle Bv_{1},w_{1}\rangle,\ldots,\langle Bv_{n},w_{n}\rangle)^{\prime}$.
@@ 2.2 Cointegrated $\mathcal{H}$-valued time series
Subsequently, we will consider cointegrated time series introduced in this section, but some additional restrictions will be imposed for our asymptotic analysis in Section 3.
@@ 3 Proposed model
Let $\{x_{t}\}_{t\geq 1}$ be a cointegrated $\mathcal{H}$-valued time series, as detailed in Section 2.2, which induces a bipartite partition of $\mathcal{H}$ into a nonstationary subspace $\mathcal{H}^{N}$ and a stationary subspace $\mathcal{H}^{S}$. We consider the following data-generating mechanism: for $f:\mathcal{H}\to{\mathcal{H}}_{y}$,
$y_{t}=f(x_{t})+u_{t},\quad\,\,u_{t}^{N}={\rm P}^{N}\Delta x_{t}=\sum_{j=0}^{\infty}\psi_{j}^{N}\epsilon_{t-j},\quad\text{and}\quad u_{t}^{S}={\rm P}^{S}x_{t}=\sum_{j=0}^{\infty}\psi_{j}^{S}\epsilon_{t-j}.$ (2)
@@
Throughout this paper, we assume that $x_{t}$ cannot be directly observed but that only $\tilde{x}_{t}$, observed with measurement errors, is available. As highlighted by Seong and Seo, (2025), this assumption is empirically relevant because functional observations used in practice are often incompletely observed, with only finitely many discrete realizations available to practitioners. Consequently, it is common to construct a functional observation $z_{t}$ in advance by smoothing its $n$ discrete data points $z_{t}(s_{1}),\ldots,z_{t}(s_{n})$, with $s_{j}$ included in the entire interval $a_{1},a_{2}]$, before computing estimators or test statistics. While one may disregard measurement errors for simplicity if $n$ is large enough and the data points are densely observed over $a_{1},a_{2}]$, this is often not the case in practice (our empirical application in Section 5 is an example). We develop theoretical results under the presence of measurement errors in functional variables, while also discussing how these results simplify in the absence of such errors. Accordingly, the subsequent theoretical developments remain applicable to the error-free case, which has been more commonly considered in the functional data analysis literature.
@@
where $e_{t}$ may generally be correlated with the variables $u_{t}^{N}$ and $u_{t}^{S}$ appearing (2) and (2), and it may also be serially correlated. A key assumption which we employ for our asymptotic analysis is that $e_{t-\kappa}$ (and also $e_{t+\kappa}$) for some finite $\kappa>0$ has asymptotically negligible sample (cross-)covariances with $u_{t}^{N}$, $u_{t}^{S}$, and $e_{t}$. Given that $e_{t-\kappa}$ or $e_{t+\kappa}$ is (primarily) smoothing error or random disturbance associated with a variable observed at a different time, this assumption is practically reasonable and likely satisfied even for small positive $\kappa$. Similarly, $y_{t}$ can suffer from error contamination, but as in conventional multivariate regression models, its measurement error is absorbed into $u_{t}$. This only changes the interpretation of $u_{t}$ in the subsequent analysis. In the presence of measurement errors $e_{t}$, we may rewrite (2) as follows:
@@
    (Error-contaminated case) $\kappa\geq 1$ and $\mathbb{E}[e_{t}\otimes z_{t+\ell}]=0$ for any $|\ell|\geq\kappa$, where $z_{t}=e_{t}$, $u_{t}^{N}$ and $u_{t}^{S}$; moreover, $\{e_{t}\}_{t\geq 1}$ is stationary and geometrically strongly mixing, $\mathbb{E}[\|e_{t}\|^{4}]<\infty$, $T^{-1}\sum_{t=1}^{T}e_{t}=O_{p}(T^{-1/2})$, and $T^{-1}\sum_{t=1}^{T}e_{t}\otimes z_{t+\ell}=\mathbb{E}[e_{t}\otimes z_{t+\ell}]+O_{p}(T^{-1/2})$ for any $|\ell|\geq\kappa$.
@@
Comments on Assumptions 1 and E are in order. We assume that $\mathcal{H}_{y}=\mathbb{R}$ (resp. $\mathcal{H}_{y}=\mathcal{H}$) if $y_{t}$ is scalar-valued (resp. function-valued). In the function-valued case, $y_{t}$ may be defined not on $a_{1},a_{2}]$, as $x_{t}$ is, but on a different interval, say $b_{1},b_{2}]$; extending the subsequent theoretical results to this case is straightforward, and hence assuming $\mathcal{H}_{y}=\mathcal{H}$ entails no loss of generality. In Assumption 1(a), we assume that $d_{N}$ is finite. This condition has been widely employed in the literature on nonstationary functional time series and seems empirically relevant (Chang et al., 2016b, ; Nielsen et al.,, 2023; Seo and Shang,, 2024). A wide class of functional time series satisfies this condition (see Remark 3.2). For convenience, we assume that $d_{N}$ is known, even though it is not the case in most empirical applications. However, replacing $d_{N}$ with a consistent estimator does not affect the asymptotic results to be developed. Moreover, we show in Section C.2 of the Supplementary Material that the variance-ratio testing procedure of Nielsen et al., (2023) can be used in our setting, allowing for measurement errors. Assumption 1(b) is employed to facilitate our theoretical analysis based on useful limit theorems in the existing literature (see, e.g., Bosq,, 2000). Given Assumption 1(b), Assumption 1(c) does not appear restrictive, and some primitive sufficient conditions can be found in Bosq, (2000, Chapter 2). Assumption 1(d) is a technical condition required for our asymptotic analysis. Similar assumptions were employed by Seo, (2024) in the study of the functional principal component analysis (FPCA) for cointegrated functional time series. Sufficient, but not restrictive, conditions for the weak convergence results stated in Assumption 1(d) can be found in, e.g., Berkes et al., (2013) and Seo, (2024).
@@
Assumption E states requirements on the measurement errors. Although our primary focus is on the error-contaminated case (Assumption E(a)), we will also show how the theoretical results simplify in the error-free case (Assumption E(b)) when applying the standard covariance-based approach (see Remark 3.1). If $e_{t}$ is serially independent and also independent of $u_{s}^{S}$ and $u_{s}^{N}$ for every $s$ and $t$, then, noting that (i) $\mathbb{E}[e_{t}\otimes u_{s}^{N}]=\mathbb{E}[e_{t}\otimes u_{s}^{S}]=0$ for all $s$ and $t$ in the considered scenario and (ii) $e_{t}\otimes z_{t}$ is a Hilbert-valued random variable (see Theorems 2.7 and 2.16 of Bosq,, 2000), the conditions in Assumption E(a) are satisfied under mild conditions. However, Assumption E(a) is not restricted to such a case and allows more general cases; specifically, $e_{t}$ is assumed to be uncorrelated with $z_{t+\ell}$ if $|\ell|$ is sufficiently large, while no restriction is imposed on $\mathbb{E}[e_{t}\otimes z_{t+\ell}]$ if $|\ell|$ is small. That is, for our theoretical investigation of the proposed method, we only require each of $e_{t}$, $u_{t}^{N}$, and $u_{t}^{S}$ to be uncorrelated with non-adjacent past or future measurement errors $e_{s}$. This not only appears to be a mild assumption but is also reasonable for most empirical applications.
@@
Practitioners may find it useful to examine the theoretical results for the standard FPCA-based estimator (corresponding to the case with $\kappa=0$, as will be shown) in the absence of measurement errors, since functional data may sometimes be observed accurately. To the authors’ knowledge, even in this simplified setting, no statistical theory has been established for nonstationary $y_{t}$ and $x_{t}$ (although Chang et al.,, 2024 studied the nonstationary functional AR(1) model). Accordingly, the subsequent results for $\kappa=0$ and the error-free case are also novel, motivating our explicit consideration of this scenario.
@@
We introduce additional notation. When these quantities are well-defined, we let $\bm{\mathsf{\lambda}}_{j}[A]$ be the $j$-th largest eigenvalue of a compact operator $A$, $\bm{\mathsf{v}}_{j}[A]$ be the corresponding eigenvector, and $\bm{{\Pi}}_{j}[A]$ be the orthogonal projection onto $\operatorname{span}\{\bm{\mathsf{v}}_{j}[A]\}$; that is,